A rectangle has a 200 meter perimeter, and its length is 8 times its width. What are the dimensions of the rectangle?
Answer by chriswen(106)

Let x m be the width of the rectangle.
Let 8x m be the length of the rectangle.
...
P=2(l+w)
200=2(8x+x)
100=9x
x=100/9
x=11.1
8x=88.9
Therefore the dimensions of the rectangle are 11.1 m by 88.9 m.